OpenStudy (anonymous):

can u define V=ttr^3/3

OpenStudy (mathmath333):

is this maths or physics

OpenStudy (mathmath333):

a city stores fresh water in a spherical tank. the tank has a radius of 32 feet. if they drain 140000 gal how much is left answer is 886763gal volum+4tt r^3/3 V=4tt *32^3/3 V=137258.27743 cubic ft 1ft^3=7.48052gallons 7.48.52*137258.27743-140000= 886763 gallons

OpenStudy (mathmath333):

as the tank is spherical u need to use the formula for volume of sphere.

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} V_{sphere}=\dfrac{4}{3}\pi r^3\\~\\ \end{align}}\)

OpenStudy (mathmath333):

here \(\large \color{black}{\begin{align} \pi=3.14\\~\\ r=radius \end{align}}\)

OpenStudy (anonymous):

so how do u do radius3

OpenStudy (anonymous):

(radius)3

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} r=radius=32^3=32\times 32\times 32=32768 \end{align}}\) see if that makes sense

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so 32768x3.14x.75

OpenStudy (anonymous):

no i mean1.33

OpenStudy (mathmath333):

so when tank is full total water will be \(\large \color{black}{\begin{align} V_{sphere}& =\dfrac{4}{3}\pi r^3\\~\\ & =\dfrac{4}{3}\times 3.14\times32768\\~\\ & =1.33\times 3.14\times32768\\~\\ \end{align}}\)

OpenStudy (anonymous):

136845.7216

OpenStudy (mathmath333):

http://www.wolframalpha.com/input/?i=%3D4%2F3times+%5Cpi+%5Ctimes32768%5C so \(\large \color{black}{\begin{align} V_{sphere}=137258.2 \end{align}}\)

OpenStudy (mathmath333):

a small difference happens but its ok

OpenStudy (anonymous):

how do i get 137258

OpenStudy (mathmath333):

use calculator

OpenStudy (anonymous):

i did it keeps coming up 136

OpenStudy (anonymous):

136845.7216

OpenStudy (mathmath333):

dont put 1.333 multiply with 4 and divide by 3 also use \(\Large \pi\) symbol instead of \(3.14\)

OpenStudy (mathmath333):

shall we proceed

OpenStudy (anonymous):

i dont understand X with 4 and div by 3 i understand tt though could u show it

OpenStudy (anonymous):

and what does ^ mean

OpenStudy (mathmath333):

value of \(4/3\) is not \(1.33\) its \(1.3333333333333333333333333333333.................\) so thats why we use that see the link it is of calculator http://www.wolframalpha.com/input/?i=%3D4%2F3times+%5Cpi+%5Ctimes32768%5C

OpenStudy (anonymous):

no i have that i only put 1.33

OpenStudy (mathmath333):

see this http://www.wolframalpha.com/input/?i=4%2F3%3D

OpenStudy (anonymous):

my next answer is 1ft ^3= 7.48052 gal

OpenStudy (mathmath333):

yes

OpenStudy (anonymous):

can u show the process 1ft^3

OpenStudy (anonymous):

i dont know what ^ is

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} V_{sphere}=137258.2 ~~cubic~~ feet \end{align}}\) we have to change it in gallon \(\large \color{black}{\begin{align} 1~~cubic~~ft=7.48~~gallon \end{align}}\) so \(\large \color{black}{\begin{align} V_{sphere}=7.48\times 137258.2 ~~gallon \end{align}}\)

OpenStudy (anonymous):

my formulas sheet sorry i feel stupid right now

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} V_{sphere}=7.48\times 137258.2=1026691 ~~gallon \end{align}}\)

OpenStudy (mathmath333):

when 140000 gallon id drain , total water left is \(\large \color{black}{\begin{align} water~~left=1026691-140000=886691~~ gallon \end{align}}\)

OpenStudy (anonymous):

ok so my next one is an order of 12 balls packek for shipment each ball is 9 inches in dia and weighs 14 oz if the shipment is based on vol so dia is 9 so the r+4.5

OpenStudy (anonymous):

thanks for ur help

OpenStudy (anonymous):

so what does vol= 4/3 pi x r^3

OpenStudy (mathmath333):

ball is sized as sphere so u will have to use sphere formula

OpenStudy (anonymous):

1.33 x 3.14 x r^3 means 4.5x4.5x4.5 =380.556225

OpenStudy (anonymous):

right

OpenStudy (mathmath333):

no only \(r^3=4.5^3=4.5\times 4.5\times 4.5\) \(1.33 \times 3.14 \times r^3 \\=1.33 \times 3.14\times 4.5\times 4.5\times 4.5\)

OpenStudy (anonymous):

I thougt thats what i just did i have 380 i dont have a pie botton

OpenStudy (anonymous):

the answer is 382 i need the pie botton

OpenStudy (mathmath333):

if u dont have \(\pi \) use \(3.14\)

OpenStudy (anonymous):

i did

OpenStudy (mathmath333):

by the way can u write ur complete question , it seems a little incomplete

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

A farmer normally mixes 1 1/2 gallons of pesticide solution with 5 ... How much pesticide solution should the farmer place in a 350 ... show more ...

OpenStudy (anonymous):

A farmer normally mixes 1 1/2 gallons of pesticide solution with 5 1/4 gallons of water for each acre of his fields. How much pesticide solution should the farmer place in a 350 gallon tank in order to treat his 45 acre field? Either 11.7 gallons 67.5 77.8 or 236. I got 77.8. 1 cubic foot = 7.48 gal

OpenStudy (mathmath333):

i thought ur question is this one ? ok so my next one is an order of 12 balls packek for shipment each ball is 9 inches in dia and weighs 14 oz if the shipment is based on vol so dia is 9 so the r+4.5

OpenStudy (anonymous):

ok hold on and ill past it out

OpenStudy (anonymous):

An order of 12 balls is being packed for shipment to New Jersey. Each ball is 9 inches in diameter and weighs 14 ounces. If the shipment charges are based on the volume of the balls, what is the volume of one of the balls? A) 85 cubic inches B) 286 C) 382 D) 3052 I am getting 381.5, not sure though I need this to be correct, thank you. Best Answer CUNY answered 5 years ago You need the following formula: V = (4/3)(3.14)(r^3) The diameter is 9 inches, which means the radius is 9/2. Let V = volume V = (4/3)(3.14)(9/2)^3 V = (4/3)(3.14)(81/4) V = 84.78, which can be rounded off to the nearest unit to become 85 cubic inches.

OpenStudy (anonymous):

i have 380 with this one its 382 on the paper i think becauce no TT

OpenStudy (anonymous):

i just loked up my s4 and i can use pie know

OpenStudy (mathmath333):

\(\large (\frac{9}{2})^3=\frac{729}{8}\) not \(81/4\)

OpenStudy (anonymous):

9/2 is 3x3x3 ?

OpenStudy (anonymous):

27

OpenStudy (mathmath333):

\((\dfrac{9}{2})^3=\dfrac{9\times 9\times 9}{2\times 2\times 2}\)

OpenStudy (anonymous):

ok i see

OpenStudy (mathmath333):

\(V = (4/3)(\pi)(9/2)^3 =381~~cubic~~inches\)

OpenStudy (anonymous):

u rock i see

OpenStudy (anonymous):

can we do one more

OpenStudy (mathmath333):

sry i have to go now but u can post more questions as u wish and tag people for help like this @jon_1 u can close this question and open a new post for each question

OpenStudy (anonymous):

thatk you so much for ur help again thank you